### v1

@@ Acknowledgements
We thank Ronny Bergmann for his insightful discussions on the parallel Douglas-Rachford methods in the context of Hadamard manifolds, and for his suggestions regarding the numerical experiments.
@@ Appendix A Proof of Proposition 20
Proof.
Let $u=\exp_{x}(-\theta\exp_{x}^{-1}y)$. Consider the geodesic triangle $\triangle(u,x,z)$. Then, there exists the corresponding comparison triangle $\triangle(\bar{u},\bar{x},\bar{z})\subset\mathbb{R}^{2}$ such that
@@
Suppose that the angles at the vertices $z$ and $\bar{z}$ are denoted by $\varsigma$ and $\varsigma^{\prime}$ respectively. Then, by Lemma 10, we have $\gamma\leq\gamma^{\prime}$. From Remark 9(i), we get
@@
From Remark 9(ii), we have
@@
Appendix B Nonexpansiveness of Reflection Operators
We will see the nonexpansiveness of the operators $R_{\lambda\phi}$ and $R_{\lambda\psi}$ (see Remark 35).
$R_{\lambda\phi}$ is nonexpansive. Indeed for $x=(x_{1},x_{2})^{T},y=(y_{1},y_{2})^{T}\in\mathcal{M}$, we have
$d^{2}(R_{\lambda\phi}(x),R_{\lambda\phi}(y))$ $=d^{2}\left(\left(x_{1},\frac{2(x_{2}+2a\lambda x_{1}^{2})}{1+2a\lambda}-x_{2}\right)^{T},\left(y_{1},\frac{2(y_{2}+2a\lambda y_{1}^{2})}{1+2a\lambda}-y_{2}\right)^{T}\right)$
$=(x_{1}-y_{1})^{2}+\left(x_{1}^{2}-y_{1}^{2}-\frac{2(x_{2}+2a\lambda x_{1}^{2})}{1+2a\lambda}+x_{2}+\frac{2(y_{2}+2a\lambda y_{1}^{2})}{1+2a\lambda}-y_{2}\right)^{2}$
$=(x_{1}-y_{1})^{2}+\left(\frac{x_{1}^{2}+2a\lambda x_{1}^{2}-2x_{2}-4a\lambda x_{1}^{2}+x_{2}+2a\lambda x_{2}}{1+2a\lambda}\right$.
$\quad\left.-\frac{y_{1}^{2}+2a\lambda y_{1}^{2}-2y_{2}-4a\lambda y_{1}^{2}+y_{2}+2a\lambda y_{2}}{1+2a\lambda}\right)^{2}$
$=(x_{1}-y_{1})^{2}+\left(\frac{(1-2a\lambda)(x_{1}^{2}-x_{2})}{1+2a\lambda}-\frac{(1-2a\lambda(y_{1}^{2}-y_{2}))}{1+2a\lambda}\right)^{2}$
$=(x_{1}-y_{1})^{2}+\left(\frac{1-2a\lambda}{1+2a\lambda}\right)^{2}\left(x_{1}^{2}-y_{1}^{2}-x_{2}+y_{2}\right)^{2}$
$\leq(x_{1}-y_{1})^{2}+\left(x_{1}^{2}-y_{1}^{2}-x_{2}+y_{2}\right)^{2}=d^{2}(x,y).$
$R_{\lambda\psi}$ is nonexpansive. Indeed for $x=(x_{1},x_{2})^{T},y=(y_{1},y_{2})^{T}\in\mathcal{M}$, we have
$d^{2}(R_{\lambda\psi}(x),R_{\lambda\psi}(y))$
$=d^{2}\left(\left(\frac{(1-2\lambda)x_{1}+4\lambda b}{1+2\lambda},x_{2}-\frac{8\lambda(x_{1}+2\lambda b)(x_{1}-b)}{(1+2\lambda)^{2}}\right)^{T},\right$.
$\quad\left.\left(\frac{(1-2\lambda)y_{1}+4\lambda b}{1+2\lambda},y_{2}-\frac{8\lambda(y_{1}+2\lambda b)(y_{1}-b)}{(1+2\lambda)^{2}}\right)^{T}\right)$
$=\left(\frac{(1-2\lambda)x_{1}+4\lambda b}{1+2\lambda}-\frac{(1-2\lambda)y_{1}+4\lambda b}{1+2\lambda}\right)^{2}+\left(\left(\frac{(1-2\lambda)x_{1}+4\lambda b}{1+2\lambda}\right)^{2}\right$.
$\quad\left.-\left(\frac{(1-2\lambda)y_{1}+4\lambda b}{1+2\lambda}\right)^{2}-x_{2}+\frac{8\lambda(x_{1}+2\lambda b)(x_{1}-b)}{(1+2\lambda)^{2}}+y_{2}-\frac{8\lambda(y_{1}+2\lambda b)(y_{1}-b)}{(1+2\lambda)^{2}}\right)^{2}$
$=\left(\frac{1-2\lambda}{1+2\lambda}\right)^{2}(x_{1}-y_{1})^{2}+\left(\left(\frac{1-2\lambda}{1+2\lambda}\right)^{2}x_{1}^{2}+\left(\frac{4\lambda b}{1+2\lambda}\right)^{2}+\frac{8\lambda b(1-2\lambda)x_{1}}{(1+2\lambda)^{2}}\right$.
$\quad\left.-\left(\frac{1-2\lambda}{1+2\lambda}\right)^{2}y_{1}^{2}-\left(\frac{4\lambda b}{1+2\lambda}\right)^{2}-\frac{8\lambda b(1-2\lambda)y_{1}}{(1+2\lambda)^{2}}+\frac{8\lambda(x_{1}+2\lambda b)(x_{1}-b)}{(1+2\lambda)^{2}}\right$.
$\quad\left.-\frac{8\lambda(y_{1}+2\lambda b)(y_{1}-b)}{(1+2\lambda)^{2}}-x_{2}+y_{2}\right)^{2}$
$=\left(\frac{1-2\lambda}{1+2\lambda}\right)^{2}(x_{1}-y_{1})^{2}+\left(\left(\frac{1-2\lambda}{1+2\lambda}\right)^{2}x_{1}^{2}+\frac{8\lambda b(1-2\lambda)x_{1}}{(1+2\lambda)^{2}}-\left(\frac{1-2\lambda}{1+2\lambda}\right)^{2}y_{1}^{2}\right$.
$\quad\left.-\frac{8\lambda b(1-2\lambda)y_{1}}{(1+2\lambda)^{2}}+\frac{8\lambda x_{1}^{2}}{(1+2\lambda)^{2}}-\frac{8\lambda b(1-2\lambda)x_{1}}{(1+2\lambda)^{2}}-\frac{16\lambda^{2}b^{2}}{(1+2\lambda)^{2}}\right$.
$\quad\left.-\frac{8\lambda y_{1}^{2}}{(1+2\lambda)^{2}}+\frac{8\lambda b(1-2\lambda)y_{1}}{(1+2\lambda)^{2}}+\frac{16\lambda^{2}b^{2}}{(1+2\lambda)^{2}}-x_{2}+y_{2}\right)^{2}$
$=\left(\frac{1-2\lambda}{1+2\lambda}\right)^{2}(x_{1}-y_{1})^{2}+\left(\left(\frac{1-2\lambda}{1+2\lambda}\right)^{2}x_{1}^{2}+\frac{8\lambda x_{1}^{2}}{(1+2\lambda)^{2}}\right$.
$\quad\left.-\left(\frac{1-2\lambda}{1+2\lambda}\right)^{2}y_{1}^{2}-\frac{8\lambda y_{1}^{2}}{(1+2\lambda)^{2}}-x_{2}+y_{2}\right)^{2}$
$=\left(\frac{1-2\lambda}{1+2\lambda}\right)^{2}(x_{1}-y_{1})^{2}+(x_{1}^{2}-y_{1}^{2}-x_{2}+y_{2})^{2}$
$\leq(x_{1}-y_{1})^{2}+(x_{1}^{2}-y_{1}^{2}-x_{2}+y_{2})^{2}=d^{2}(x,y).$

### v2

@@ Acknowledgements
We thank Ronny Bergmann for his insightful discussions on the parallel Douglas-Rachford methods in the context of Hadamard manifolds, and for his suggestions regarding the numerical experiments. The authors are grateful to the anonymous referees for careful reading of the manuscript and for various suggestions that helped to improve the paper.
@@ Appendix A Proof of Proposition 9
First, we recall some basic results on Hadamard manifolds.
Lemma 40 ([Martin1]).
Let $\mathcal{M}$ be a Hadamard manifold and let $x,y,z\in\mathcal{M}$. Then, we have the following:
  1. (i)
  2. (ii)
Lemma 41 ([Martin1]).
  1. (a)
  2. (b)
Lemma 42 ( [KhammaAOT2022]).
Let $\mathcal{M}$ be an Hadamard manifold and $x,y,z\in\mathcal{M}$. Then
$\|\exp_{x}^{-1}z-P_{x,y}\exp_{y}^{-1}z\|\leq d(x,y).$
Proof of Proposition 9. Let $u=\exp_{x}(-\theta\exp_{x}^{-1}y)$. Consider the geodesic triangle $\triangle(u,x,z)$. Then, there exists the corresponding comparison triangle $\triangle(\bar{u},\bar{x},\bar{z})\subset\mathbb{R}^{2}$ such that
@@
Suppose that the angles at the vertices $z$ and $\bar{z}$ are denoted by $\varsigma$ and $\varsigma^{\prime}$ respectively. Then, by Lemma 41, we have $\gamma\leq\gamma^{\prime}$. From Lemma 40(i), we get
@@
From Lemma 40(ii), we have
@@
Appendix B Proof of Proposition 15
For the proof of Proposition 15, we first recall [Leon2013, Proposition 4.1] for sectional curvature $k=0$.
Proposition 43 ([Leon2013]).
Let $\mathcal{M}$ be a Hadamard manifold with zero sectional curvature and let $x,y,x^{\prime},y^{\prime},a,b\in\mathcal{M}$ satisfy $x^{\prime}=\exp_{a}\bigl(-\exp_{a}^{-1}x\bigr)\quad\text{and}\quad y^{\prime}=\exp_{b}\bigl(-\exp_{b}^{-1}y\bigr).$ Suppose that $d(a,b)\leq d\bigl(\gamma(x,a;t),\gamma(y,b;t)\bigr)$ for all $t\in[0,1]$. Then $d(x^{\prime},y^{\prime})\leq d(x,y).$
Proof of Proposition 15. By the definition of the reflection, we have
$R_{\lambda f}(x)=\exp_{\operatorname*{Prox}_{\lambda f}(x)}\!\left(-\exp_{\operatorname*{Prox}_{\lambda f}(x)}^{-1}(x)\right),\quad R_{\lambda f}(y)=\exp_{\operatorname*{Prox}_{\lambda f}(y)}\!\left(-\exp_{\operatorname*{Prox}_{\lambda f}(y)}^{-1}(y)\right).$
Since $\operatorname*{Prox}_{\lambda f}$ is a firmly nonexpansive operator, it follows from (10) that
$d(\operatorname{Prox}_{\!\lambda f}(x),\operatorname{Prox}_{\!\lambda f}(y))\leq d(\gamma(x,\operatorname{Prox}_{\!\lambda f}(x);t),\gamma(y,\operatorname{Prox}_{\!\lambda f}(y);t))\text{ for all }t\in[0,1].$
Therefore, by Proposition 43, we obtain $d\bigl(R_{\lambda f}(x),R_{\lambda f}(y)\bigr)\leq d(x,y).$